mikewwe13:

What is the explicit rule for this geometric sequence? a1 =−15;an= 1/5 ⋅ an−1 A. an = −15(1/5)n−1 B. an = 15(−1/5)n−1 C. an= 1/5 ⋅ (−15)n−1 D. an = −1/5 ⋅ 15n−1

Vocaloid:

the common ratio is 1/5 and the first term is -15 write an expression where an = first term * common ratio ^ (n-1)

mikewwe13:

a1 = 1/5 x -15

mikewwe13:

but the answer is B

Vocaloid:

an = first term * common ratio ^ (n-1) replace "first term" with -15 and "common ratio" with (1/5)

mikewwe13:

an = -15 x (1/5) ^ (n - 1)

Vocaloid:

good and which answer choice matches?

mikewwe13:

A

Vocaloid:

good.

Vocaloid:

is this the exact question? the sequence doesn't appear to be geometric

mikewwe13:

WAIT

Vocaloid:

oh wait I get it, the first term is 4/3

mikewwe13:

What is the common ratio for the geometric sequence? 4/3, 4, 12, 36,... Enter your answer in the box. ___________

mikewwe13:

yes

mikewwe13:

my mistake

Vocaloid:

anyway, you want to divide the second term by the first term

Vocaloid:

4/(4/3) = ?

mikewwe13:

3

Vocaloid:

good so that's your common ratio

mikewwe13:

The function h (m) = 10 + 4m represents the height of a vine, in centimeters, m months after the vine was transplanted. What does the value 10 represent in this situation? A. The vine grows 10 cm each month. B. The vine was 10 cm tall when it was transplanted. C. The vine grew for 10 months. D. The vine was 10 months old when it was transplanted.

mikewwe13:

The answer is c, since at month 0 we have h(m) = 10 + 4*0 h(m) = 10 cm

mikewwe13:

CORRECT ?

Vocaloid:

10 is the initial value, or the value at m = 0

mikewwe13:

B

Vocaloid:

good, b.

mikewwe13:

The function p (d) = 5 ⋅ 4^d−1 represents the number of people who receive a certain e-mail on day d of an e-mail experiment. What does the value 5 represent in this situation? A. Every day, the number of people receiving the e-mail increases by 5. B. Five people received the e-mail on the first day of the experiment. C. Every day, the number of people receiving the e-mail increases by 5%. D. There are 5 people running the experiment.

Vocaloid:

hint: 5 is the initial value, what does that mean in the context of the problem?

mikewwe13:

that is the ordinary differential equation that is put together with a specified value

Vocaloid:

that is not relevant to the context of this problem, this is not a differential equation.

Vocaloid:

since the output represents the number of people receiving the email and d = number of days, the initial value is the number of people receiving the email on day 1

Vocaloid:

please stop pasting irrelevant information

Vocaloid:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Vocaloid since the output represents the number of people receiving the email and d = number of days, the initial value is the number of people receiving the email on day 1 \(\color{#0cbb34}{\text{End of Quote}}\) use this as a hint to pick the correct answer

mikewwe13:

it's B

Vocaloid:

good.

mikewwe13:

The function q (w) = 3 + 5(w − 1) represents the number of quarters in a bowl on week w. What does the value 5 represent in this situation? A. Quarters were added to the bowl for 5 weeks. B. There were 5 quarters in the bowl on Week 1. C. Five quarters are added to the bowl every week. D. The value of the quarters in the bowl on Week 1 was $5.

Vocaloid:

as a hint, 5 is the rate of change

mikewwe13:

D

Vocaloid:

5 is the rate at which we are adding quarters to the bowl

mikewwe13:

A

Vocaloid:

~rate~ not time

mikewwe13:

it's not A OR d it's either B or C C

Vocaloid:

good, C.

mikewwe13:

The function s(t) represents the value of a toy t years after it was purchased. s(t) = 65(0.87)^t What does the value 65 represent in this situation? A. The initial value of the toy is $65. B. Every year the value of the toy increases by $65. C. Every year the value of the toy decreases by $65. D. The value of the toy after t years is $65.

Vocaloid:

in the geometric sequence formula, the number on the left side being multiplied to r^(n-1) is the ~initial value~

mikewwe13:

A

Vocaloid:

good

mikewwe13:

The function c(t) represents the total cost for an artist to draw t pictures. c(t) = 34.5t + 127.5 What does the value 34.5 represent in this situation? A. For each picture drawn, the total cost decreases by $34.50. B. The initial cost is $34.50. C. For each picture drawn, the total cost increases by $34.50. D. The total cost to draw t pictures is $34.50.

Vocaloid:

since 34.5 is being multiplied by the variable t we can presume that this is the cost per painting

Vocaloid:

which answer makes the most sense given this information?

mikewwe13:

B

Vocaloid:

$34.5 is not the initial cost, it is the cost per painting

mikewwe13:

D

Vocaloid:

"cost per painting" means that each painting costs $34.50

Vocaloid:

each picture increases the cost by $34.50

mikewwe13:

it's not B or D it's either A or C C

Vocaloid:

good, C.

mikewwe13:

ok thank you for all your help, have a wonderful day Vocaloid